OpenStudy (anonymous):

Convert the function into standard form. Show your work. y=2(x-4)(x+3)

jigglypuff314 (jigglypuff314):

first foil the (x-4)(x+3) part then distribute the 2 then simplify...

OpenStudy (anonymous):

and when you are done, tell your math teachers (gently) that \[y=2(x-4)(x+3)\] is an EQUATION, not a function

OpenStudy (precal):

If you replace y with f(x), such as f(x)=2(x-4)(x+3) then I believe will make it a function.

OpenStudy (anonymous):

i suppose so, or you could write \[f:=\{(x,y)|y=2x^2-2x-24\}\] or something, but if it is supposed to be a function you have to at least write it as one

OpenStudy (anonymous):

f(x)=2(x-4)(x+3) so thats the function

jigglypuff314 (jigglypuff314):

first foil the (x-4)(x+3) part then distribute the 2 then simplify...

OpenStudy (anonymous):

so whats the function?

jigglypuff314 (jigglypuff314):

foil (x-4)(x+3) so you get 2(x^2 - 4x + 3x - 12) then distribute the 2 then add like terms to simplify (It would be against this site's code of conduct to give the actual answers...)

OpenStudy (anonymous):

I don't get it :/

OpenStudy (anonymous):

Can someone explain how to do this

OpenStudy (anonymous):

lets see y = 2 (x-4 )(x+3 ) y= 2 ( x (x+3 ) -4 (x+3 ) y= 2 ( x pow (2) + 3x - 4x -12 ) y = 2 ( x pow (2) -x -12 ) y = 2x pow (2) -2x -24

Directrix (directrix):

y = 2(x-4)(x+3) y = 2(x^2 - 4x + 3x - 12) y = 2( x^2 - x - 12) y = 2x² - 2x - 24 All that comes from above. This is what is left to do: Convert the function into standard form. Look in your text to determine how standard form form of a quadratic equation is defined. Some texts use standard form and vertex form to mean the same. Others do not.

OpenStudy (anonymous):

@ directrix can u help me how u gave power above x to be done here ?

Directrix (directrix):

@vinurules I copied and pasted the 2 onto the x to get x². You can copy this and save it for your use. ² Many people here use LaTex for math symbols. There is a section on how to use that here: http://openstudy.com/study#/groups/LaTeX%20Practicing!%20%3A)

Directrix (directrix):

y = 2(x-4)(x+3) y = 2(x^2 - 4x + 3x - 12) y = 2( x^2 - x - 12) y = 2x² - 2x - 24 All that comes from above. This is what is left to do: Convert the function into standard form. See attachment.

Directrix (directrix):

^^ That is also the vertex form for the equation of a parabola where (h,k) is the vertex of the parabola;n a ≠ 0. @ShantelT1

Directrix (directrix):

For reference, attached is the standard form of a quadratic equation in standard form.

OpenStudy (anonymous):

So this is the function?

Directrix (directrix):

The line in you posted question is this: Convert the function into standard form. So, I assume that the task is to write the given function in the form as described in the attachment below.

Directrix (directrix):

That means there is more work to do.

OpenStudy (anonymous):

F(X)a(x-h)^2

OpenStudy (anonymous):

thats what i ment

Directrix (directrix):

y = 2(x-4)(x+3) y = 2(x^2 - 4x + 3x - 12) y = 2( x^2 - x - 12) Complete the square: y = 2 ( x ^2 - x + 1/4) - 24 - 1/2 and continue.

OpenStudy (precal):

directrix did a great job explaining